OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 1 less than or equal to 12 or 2x greater than or equal to 30 A number line with an open circle on 13, shading to the left, and an open circle on 15, shading to the right. A number line with a closed circle on 13, shading to the left, and a closed circle on 15, shading to the right. A number line with an open circle on 13, an open circle on 15, and shading in between. A number line with a closed circle on 13, a closed circle on 15, and shading in between.

OpenStudy (anonymous):

@Preetha @Mertsj @hartnn ?

OpenStudy (anonymous):

@Hero @satellite73

OpenStudy (mertsj):

Solve each part.

OpenStudy (mertsj):

\[x-1\le 12\]

OpenStudy (mertsj):

Add 1 to both sides.

OpenStudy (mertsj):

\[2x \ge 30\]

OpenStudy (mertsj):

Divide both sides by 2

OpenStudy (anonymous):

Ohhh, alright. I get it. Thank you.

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

Which one is it?

OpenStudy (anonymous):

Working it out right now.

OpenStudy (anonymous):

\[x \le 13 \] and \[x \ge 15\]

OpenStudy (anonymous):

@Mertsj ?

OpenStudy (mertsj):

Perfect. Good for you!!!

OpenStudy (anonymous):

Thank you!

OpenStudy (mertsj):

yw